A company’s CEO wanted to estimate the percentage of defective product per shipment. In a sample containing 400 products, he found 25 defective products.

(a) Find a 95% confidence interval for the true proportion of defective product. Show your calculations and/or explain the process used to obtain the interval.
(b) Interpret this confidence interval and write a sentence that explains it.

Answer by stanbon(75887)

A company’s CEO wanted to estimate the percentage of defective product per shipment. In a sample containing 400 products, he found 25 defective products.
-------
(a) Find a 95% confidence interval for the true proportion of defective product. Show your calculations and/or explain the process used to obtain
the interval.
p-hat = 25/400 = 0.0625
----
ME = 1.96*sqrt[0.0625*0.9375/400] = 0.0237
95%CI: 0.0625-0.0237 < p < 0.0625+0.0237
95%CI: 0.0388 < p < 0.0862
============================================
(b) Interpret this confidence interval and write a sentence that explains it.
With 95% confidence we can say that the proportion of defective products
is between 3.88% and 8.62%
================================
Cheers,
Stan H.
================================